OpenStudy (anonymous):

show me how to work 4y=12y^2

OpenStudy (anonymous):

what have you done so far?

OpenStudy (amistre64):

-4y from each side; then factor out a y and see which y values make it equal zero

OpenStudy (anonymous):

did you just say subtract 4y from both sides?

OpenStudy (anonymous):

nothing i dont understand how to work it

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

umm thats wrong.... more like divide.

OpenStudy (anonymous):

thats what i thought it would be division

OpenStudy (amistre64):

4y = 12y^2 ; -4y 0 = 12y^2 -4y ; factor out 4y 0 = 4y(3y - 1) ; y = 0 or 1/3

OpenStudy (anonymous):

^2 means squared right that's why i typed it like that because 12y squared

OpenStudy (anonymous):

so 4y=12y squared

OpenStudy (amistre64):

am i still 'wrong' ? lol

OpenStudy (anonymous):

oh sorry yeah do it like he said and check using quadratic.

OpenStudy (anonymous):

ok i'm a 9th grader so i'm a bit confused lol

OpenStudy (amistre64):

you can also /4y if you want 1 = 3y ; y=1/3 but your missing something because... 4y CAN be a zero, which this way does not allow for

OpenStudy (amistre64):

if it can be zero; stay away from division

OpenStudy (anonymous):

\[4y=12y^2\]\[4y-4y=4y^2-4y\]\[0=4y^2-4y\]\[4(y^2-y)=0\]\[4(y(y+1))=0\]

OpenStudy (anonymous):

yes good thinking ^-^

OpenStudy (anonymous):

*12y^2

OpenStudy (amistre64):

when you know it can never be zero; then division is fine ;)

OpenStudy (anonymous):

ok ty